1.      What is the value of x after each of these statements is encountered in a computer program, if x = 1 before the statement is reached?

a)      if x + 2 = 3 then x := x + 1


1
Expert's answer
2022-04-26T13:28:13-0400

Since the statement "1+2=3" is true, it follows that after the command "if x + 2 = 3 then x := x + 1" the value of "x" is equal to "1+1=2."